OpenStudy (anonymous):

Solve the following system of linear equations using any algebraic method. If possible, check your solution. 8x-6y=14 12x-9y=18

OpenStudy (anonymous):

please help.. thanks..

OpenStudy (anonymous):

i think these two equation are just parallel to each other.

OpenStudy (anonymous):

yes i think that is true but i need to solve it..

OpenStudy (anonymous):

cuz its telling me to show all the work you do..

OpenStudy (anonymous):

well when the equation are parallel to each other there is no solution and you can't get the value of x and y cuz they never intersect each other

OpenStudy (anonymous):

ok thanks thats weird though, its from a quiz that i have to turn in to the school...i guess i cant solve it.

OpenStudy (anonymous):

actually the equation look like 2(4x-3y)=12 and in next step u'll get 4x-3y=7 and again for the next equation you'll get the same like 4x-3y=6

OpenStudy (anonymous):

so it doesn't have any solution